If 20 ladies lost 29 gloves - some lost one glove and some lost both gloves. how many lost one glove ?
Answer by richwmiller(17219)

a+b=20,
2a+b=29
a = 9, b = 11
9 lost both (18) and 11 lost one
18+11=29